# Misiurewicz points and subhyperbolicity in unicritical algebraic correspondences

### v1

@@ 5. Proof of Theorem B
By definition, every branch is holomorphic and defined on a region, that is, a nonempty, open and connected subset of the plane (see page 3.4). Every branch $f$ of $\mathbf{g}$ determines a connected set
@@
which is contained in some component of the sheaf space. By Theorem A, this component must be some $\mathfrak{S}(\mathbf{g}_{\alpha})$. Hence $f$ is a branch of $\mathbf{g}_{\alpha}.$ From Theorem A we know that every germ $\mathfrak{f}$ in $\mathfrak{S}(\mathbf{g}_{\alpha})$ can be continued along any curve in $\mathbb{D}$ starting at $\pi(\mathfrak{f}).$ It follows from Corollary 3.1 that $f$ can be extended to a branch $F$ defined on $\mathbb{D}.$
@@ 6. Uniform contraction on compact sets
There exists a conformal disk $V$ containing $a_{j}$ such that every connected component of $\varphi^{-1}(V)$ projects onto $V$ by means of a proper map. The group of deck transformations acts transitively on such components. Since the set o branch points is locally finite, by reducing $V$ if necessary we may assume that each component contains only on branch point, which must be necessarily projected to $a_{j}$ by $\varphi.$ Let $\mathcal{E}$ denote the unique component of $\varphi^{-1}(V)$ containing $x_{j}.$ Then $\varphi:\mathcal{E}{\setminus}\{x_{j}\}\to V{\setminus}\{a_{j}\}$ is $d$-fold covering map, where $d=\nu_{a}(a_{j}).$
The set $\tilde{\mathcal{D}}_{j}$ in the statement is constructed as follows. Fix any point $z_{0}$ in $V{\setminus}\{a_{j}\}$. Any $z$ in $\mathcal{D}{\setminus}\{a_{j}\}$ can be joined to $z_{0}$ by a continuous curve $\gamma$ in $\mathcal{D}{\setminus}\{a_{j}\}$ starting at $z_{0}.$ The analytic continuation of a germ of $(\varphi|_{\mathcal{E}})^{-1}$ at $z_{0}$ can be continued along $\gamma$ and leads to a germ at $z$ which is represented by $(g_{1},z)$, where $g_{1}$ is branch of $\varphi^{-1}$ locally defined at $z.$ For every $z$ in $\mathcal{D}{\setminus}\{a_{j}\},$ we let $\mathbf{h}(z)$ denote the set of all $g_{1}(z)$ such that $(g_{1},z)$ is obtained by analytic continuation of a germ of $(\varphi|_{\mathcal{E}})^{-1}$ at $z_{0}$ along a curve $\gamma$ in $\mathcal{D}{\setminus}\{a_{j}\}$ joining $z_{0}$ to $z$, as described previously. We set $\mathbf{h}(a_{j})=x_{j}.$ By the Monodromy Theorem, $\mathbf{h}(z)$ does not depend on the initial choice of $z_{0}$ and $\mathbf{h}(V)=\mathcal{E}.$ The set $\mathcal{C}=\mathbf{h}(\mathcal{D})$ – which is constructed using analytic continuations along curves – is naturally path connected and satisfies
@@
In order to show that $\overline{\mathcal{C}}\subset\mathbb{D}$, assume that $\mathcal{D}=g(D)$, where $g$ is a univalent map defined on a region containing the closure of a round disk $D.$ For a small $\epsilon>0$, the $\epsilon$-neighborhood of $D$ is contained in the domain of $g.$ Let $\mathcal{D}_{\epsilon}=g(D_{\epsilon}).$ Using the same process in the construction of $\mathcal{C},$ we construct another region $\mathcal{C}_{1}$ containing the closure of $\mathcal{C}$ by considering all possible analytic continuations of germs of $(\varphi|_{\mathcal{E}})^{-1}$ along curves in $\mathcal{D}_{\epsilon}$. It should be noticed, however, that now $\mathcal{D}_{\epsilon}$ might contain a finite set of ramified points, and as a consequence, some analytic continuations do not lead to a germ, but rather to a branch point of $\varphi$ which we add to $\mathcal{C}_{1}$ in the process of its construction. It follows that $\overline{\mathcal{C}}\subset\mathcal{C}_{1}\subset\mathbb{D}.$
If $K$ is a compact subset of $\mathcal{D}_{j},$ then by (6.1) the intersection $\varphi^{-1}(K)\cap\partial\mathcal{C}$ is empty. Hence
@@
is compact. Since $K$ is arbitrary, $\varphi:\mathcal{C}\to\mathcal{D}$ is proper, with the same degree $d$ of $\varphi|_{\mathcal{E}}.$ Finally, we will prove that $\mathcal{C}=\tilde{\mathcal{D}}_{j}.$ Since $\mathcal{C}$ is connected and $x_{j}\in\mathcal{C}$, it follows that $\mathcal{C}\subset\tilde{\mathcal{D}}_{j}.$ For the other inclusion, we consider a curve $\zeta$ in $\tilde{\mathcal{D}}_{j}$ joining some $w_{0}\in\mathcal{E}$ to an arbitrary $w$ in $\tilde{\mathcal{D}}_{j}.$ Without loss of generality, we may assume that $\gamma=\varphi\circ\zeta$ contains no ramified points; by Theorem 3.2, $\zeta$ is given by an analytic continuation continuation of a germ of $(\varphi|_{\mathcal{E}})^{-1}$ along $\gamma$, thereby showing that $w$ belongs to $\mathcal{C}.$
@@
for every $G$ in $\mathcal{G}_{1}$ and every $w\in{\tilde{\mathcal{D}}}_{1}.$ In Diagram (6.2), $\varphi$ is an isometry on each disk (using the hyperbolic metric $\mu$ on the unit disk and the orbifold metric $\rho$ on $\mathbb{C}{\setminus}R$). We conclude that $\|g^{\prime}(z)\|_{\rho}<\eta_{1},$ whenever $z\in V_{1}$, for every univalent branch $g$ defined on $V_{1}.$ We may take $V_{1}$ as the open ball $B(\zeta_{0},r_{0})$ described in the statement. The Lemma follows in the case $a_{j}=a.$ The general case involves a similar reasoning (using another sequence of univalent maps just like Diagram (6.2), with the help of Lemma 6.1). ∎
@@
Follows from Theorem B and Lemma 6.2, using a standard compactness argument, with a finite covering by relatively compact open sets on which all (finitely many) branches of $\mathbf{f}_{a}^{-1}$ contracts the orbifold metric by a uniform factor. Lemma 6.2 is applied to find this uniform factor on a neighborhood of every ramified point. Using the polar representation of $\mathbf{f}_{a}^{-1}$ and the fact that $p/q>1$ it is possible to show that $g$ extends to a univalent branch on $B(\zeta_{0},r_{0})$ by computing $g$ explicitly using polar coordinates. ∎
7. Sub-hyperbolicity
If $a$ is a Misiurewicz point, then by definition the critical point has a unique bounded forward orbit, the pre-periodic critical orbit of $\mathbf{f}_{a}$. This critical orbit is precisely the set of ramified points which are in $K_{a}$. (Recall that every ramified point is a singularity of the orbifold metric). Since both $K_{a}$ and $\mathbb{C}{\setminus}R$ are backward invariant, it follows that $K_{a}{\setminus}R$ is also backward invariant under $\mathbf{f}_{a}.$ The following result guarantees the uniform expansion of the orbifold metric within a neighborhood of $K_{a}.$
Theorem 7.1 (Sub-hyperbolicity).
Suppose that $a$ is a Misiurewicz point for the family (2.1). Then $\mathbf{f}_{a}$ expands the orbifold metric $\rho$ on a neighborhood of $K_{a}$ by a uniform factor. More precisely, there exist an open set $V$ containing $K_{a}$ and a constant $\eta\in(0,1)$ such that $R\cap V$ is the pre-periodic critical orbit of $\mathbf{f}_{a}$, and for every univalent branch $g$ of $\mathbf{f}_{a}^{-1}$ defined on a region $W\subset V,$ we have $\|g^{\prime}(w)\|_{\rho}<\eta$, for every $w\in W{\setminus}R.$
@@ 7. Sub-hyperbolicity
Since $R$ is locally finite and $K_{a}$ is compact, there exists an $\epsilon$-neighborhood $V=(K_{a})_{\epsilon}$ such that $R\cap V=R\cap K_{a}.$
@@
For every ramified point $a_{j}$ of the canonical orbifold associated to a Misiurewicz point there exists a coordinate chart $z=\phi(\zeta)$ defined for $\zeta$ in neighborhood of $a_{j}$, with $\phi(a_{j})=0$, such that the expression of orbifold metric with respect to the local uniformizing parameter $z$ around zero becomes
@@
where $d=\nu(a_{j})$ and $\rho_{j}$ is a $C^{\infty}$ and strictly positive function defined on a round disk centered at 0 which satisfy $\rho_{j}(e^{2\pi i/d}z)=\rho_{j}(z),$ for any $z$ in the domain of $\rho_{j}.$
@@
The statement can be generalized for any regular branched covering $\varphi$ (not necessarily the one used in this paper). Let $x_{j}$ be a branch point. Up to compositions with translations, $\varphi$ is locally conjugate to $z^{\nu(a_{j})}$, where $a_{j}=\varphi(x_{j})$, as a consequence of Böttcher’s Theorem. The expression (7.1) is found in [Milnor, p. 210], using the branched covering $z^{\nu(a_{j})}$ around zero, computing directly the push-forward of the metric. ∎
Recall that in the case of a Misiurewicz point, the pre-periodic critical orbit of $\mathbf{f}_{a}$ contains a unique cycle $\alpha(a).$
@@
As a bounded orbit, every cycle is in the filled Julia set. The points of $\alpha(a)$ are in $R$ and none of them is the critical point. Denote the points of $\alpha(a)$ by
@@
There exist univalent branches $f_{i}$ of $\mathbf{f}_{a}$, each of which locally defined at $\zeta_{i-1}$, with $f_{i}(\zeta_{i-1})=\zeta_{i}$. Consider the composition
@@
Let $K=K_{a}.$ Recall that the points of the pre-periodic orbit of zero coincides with $R\cap K,$ where $R$ is the set of ramified points. Let $R^{*}=R{\setminus}\{0\}.$ We know that every point $z$ of $K\cap R^{*}$ has only one image in $K$ (otherwise there would be another bounded forward orbit of zero) and the other q1q-1 points of the image of $z$ are outside of the compact set $K_{a}.$ By continuity, the correspondence maps a small conformal disk containing $z\in K\cap R^{*}$ to $q$ disjoint conformal disks, and only one of them intersects $K.$ This determines at each such $z$ an especial univalent branch $g_{z}$ of the correspondence that sends a conformal disk containing $z$ to the aforementioned conformal disk intersecting $K.$ The disjoint union of such small disks determines a neighborhood $V_{\alpha}$ of $K\cap R^{*}$ such that if the correspondence sends a point $\zeta\in V_{\alpha}$ to another point $\zeta^{\prime}$ in $V_{\alpha}$, then $\zeta^{\prime}=g_{z}(\zeta)$, for some $z\in K\cap R^{*}.$ This creates on $V_{\alpha}$ a especial regime of iteration; and since $\alpha(a)$ is repelling, any infinite forward orbit contained in $V_{\alpha}$ is pre-periodic and eventually coincides with $\alpha(a).$
@@
For any conformal disk $U$ intersecting $K_{a}$, there exists a finite sequence of nonconstant holomorphic maps $(f_{j})_{0}^{n}$ defined on $U$ such that $f_{j}$ is a branch of $\mathbf{f}_{a}^{j},$ $f_{n}(U)$ contains zero and $f_{j}^{\prime}(z)\neq 0$ whenever $z\in U$ and $j\in[0,n]\cap\mathbb{Z}.$ Since $U$ is arbitrary, (7.4) follows.
@@
Fix a bounded forward orbit $(z_{i})_{0}^{\infty}$ of a point in $U\cap K_{a}.$ If $0\in U$ or some point of this orbit is a critical point, then there is nothing to prove. Otherwise, by the Monodromy Theorem and its Corollary 3.1 there exists a unique branch $f_{1}$ of $\mathbf{f}_{a}$ defined on $U$ which sends $z_{0}$ to $z_{1}$. If $f_{1}(U)$ contains 0, then there is nothing to prove. Otherwise, we may consider the holomorphic multifunction $\mathbf{f}_{a}\circ f_{1}$ defined on $U$, and using analytic continuation as before, we find a unique branch $f_{2}$ of $\mathbf{f}_{a}\circ f_{1}$ defined on $U$ which sends $z_{0}$ to $z_{2}.$ Unless $f_{2}(U)$ contains the critical point, $f_{3}$ is by definition the unique branch of $\mathbf{f}_{a}\circ f_{2}$ defined on $U$ which sends $z_{0}$ to $z_{3}.$ If we are fortunate, this argument terminates at a branch $f_{n}$ of $\mathbf{f}_{a}^{n}$ defined over $U$ such that $f_{n}(U)$ contains zero, and there is nothing else to prove. Otherwise, we find a sequence of maps $f_{n}:U\to\{0,a\}$ which is normal, by Montel’s Theorem. However, we will see that $f_{n}$ can never be a normal family. Suppose first that the set of subsequential limits of $(z_{i})_{0}^{\infty}$ is contained in $R\cap K.$ Then all but finitely many $z_{n}$ belong the the neighborhood $V_{\alpha}$ of $K\cap R^{*}$. Hence $(z_{i})_{0}^{\infty}$ eventually coincides with the repelling cycle $\alpha,$ which implies $f_{n}^{\prime}(z_{0})\to\infty$, as a consequence of the chain rule. We conclude from the Weierstrass convergence theorem that $f_{n}$ is not normal. If the set of subsequential limits of the bounded orbit $(z_{i})_{0}^{\infty}$ is not contained in $K\cap R,$ then a subsequence $z_{n_{k}}$ converges to some $w_{0}$ in $K{\setminus}R.$ Suppose for a moment that $f_{n}$ is a normal family. Since $f_{n_{k}}(z_{0})$ converges to $w_{0},$ no subsequence of $f_{n_{k}}$ escapes to infinity. By normality, we may replace $f_{n_{k}}$ by one of its convergent subsequences, so that $f_{n_{k}}$ converges locally uniformly to some holomorphic function $g$ defined on $U$, as well as $f^{\prime}_{n_{k}}$ converges locally uniformly to $g^{\prime}.$ Since $R$ is forward invariant and $K\cap R$ is the unique bounded forward orbit of zero, it is possible to show that if one point of the sequence $(z_{i})_{0}^{\infty}$ enters $R$, then subsequent terms never leave $R$ and are eventually trapped in the repelling cycle $\alpha(a)$, a possibility that is ruled out by the fact $z_{n_{k}}\to w_{0}\notin R.$ Hence we are allowed to evaluate the norm of $f_{n_{k}}^{\prime}$ with respect to the orbifold metric $\rho$, concluding from Theorem 7.1 and the chain rule that $\|f_{n_{k}}^{\prime}\|_{\rho}$ explodes to infinity as $k\to\infty$; nevertheless,
@@
As described in (7.2) and (7.3), the composition of the univalent branches along the points $\zeta_{j}$ of the repelling cycle $\alpha$ yields a holomorphic $f$ map defined on neighborhood of $\zeta_{0}$ with a repelling fixed point at $\zeta_{0}=f(\zeta_{0}).$ The inverse $f^{-1}$ maps a conformal disk $\Omega$ around $\zeta_{0}$ into itself, with a geometrically attracting fixed point of multiplier $\lambda_{0}$ at $\zeta_{0}.$ Let $\mathbb{D}_{r}$ denote the open disk of radius $r$ centered at zero. Using the Königs linearization theorem, we may assume that $\Omega=\varphi(\mathbb{D}_{r}),$ where $\varphi$ is a conformal conjugacy between $z\mapsto\lambda_{0}z$ and $f^{-1}$.
By Lemma 7.2, $f_{n}(U)=V$ is a neighborhood of zero, for then $f_{n}(w_{0})=0.$ Then $W=\mathbf{f}_{a}(V)$ is a connected neighborhood of $a.$
There exists a conformal map $h$ defined on a neighborhood of $a$ which is a branch of some iterate $\mathbf{f}_{a}^{k_{0}}$ and sends $a$ to $\zeta_{0}.$ By choosing $s\in(0,r)$ sufficiently small, we may assume that $h$ maps a neighborhood $W_{1}$ of $a$ contained in $W$ biholomorphically onto $\Omega_{1}=\varphi(\mathbb{D}_{s})\subset\Omega.$ Let $V_{1}$ denote $\mathbf{f}_{a}^{-1}(W_{1})$, which is a connected neighborhood of zero contained in $V.$ If we reduce the size of $s$, then both sets $V_{1}$ and $W_{1}$ will shrink accondingly; since $f_{n}$ is conformal and sends $w_{0}$ to zero, by reducing $s$ if necessary we may assume that $f_{n}$ sends a small connected neighborhood $U_{2}\subset U_{1}$ of $w_{0}$ biholomorphically onto $V_{1}.$ It is important to notice that $U_{2}$ also depends on $s$, and if $s\to 0$, then $U_{2}$ will shrink to $\{w_{0}\}.$

### v2

@@ 5. Proof of Theorem B
By definition, every branch is holomorphic and defined on a region, that is, a nonempty, open and connected subset of the plane (see page 3). Every branch $f$ of $\mathbf{g}$ determines a connected set
@@
which is contained in some component of the sheaf. By Theorem A, this component must be some $\mathfrak{S}(\mathbf{g}_{\alpha})$. Hence $f$ is a branch of $\mathbf{g}_{\alpha}.$ From Theorem A we know that every germ $\mathfrak{f}$ in $\mathfrak{S}(\mathbf{g}_{\alpha})$ can be continued along any curve in $\mathbb{D}$ starting at $\pi(\mathfrak{f}).$ It follows from Corollary 3.1 that $f$ can be extended to a branch $F$ defined on $\mathbb{D}.$
@@ 6. Uniform contraction on compact sets
There exists a conformal disk $V$ containing $a_{j}$ such that every connected component of $\varphi^{-1}(V)$ projects onto $V$ by means of a proper map. The group of deck transformations acts transitively on such components. Since the set of branch points is locally finite, by reducing $V$ if necessary we may assume that each component contains only one branch point, which must be projected to $a_{j}$ by $\varphi.$ Let $\mathcal{E}$ denote the unique component of $\varphi^{-1}(V)$ containing $x_{j}.$ Then $\varphi:\mathcal{E}{\setminus}\{x_{j}\}\to V{\setminus}\{a_{j}\}$ is a $d$-fold covering map, where $d=\nu_{a}(a_{j}).$
The set $\tilde{\mathcal{D}}_{j}$ in the statement is constructed as follows. Fix any point $z_{0}$ in $V{\setminus}\{a_{j}\}$. Any $z$ in $\mathcal{D}{\setminus}\{a_{j}\}$ can be joined to $z_{0}$ by a continuous curve $\gamma$ in $\mathcal{D}{\setminus}\{a_{j}\}$ starting at $z_{0}.$ The analytic continuation of a germ of $(\varphi|_{\mathcal{E}})^{-1}$ at $z_{0}$ can be continued along $\gamma$ and leads to a germ at $z$ which is represented by $(g_{1},z)$, where $g_{1}$ is a branch of $\varphi^{-1}$ locally defined at $z.$ For every $z$ in $\mathcal{D}{\setminus}\{a_{j}\},$ we let $\mathbf{h}(z)$ denote the set of all $g_{1}(z)$ such that $(g_{1},z)$ is obtained by analytic continuation of a germ of $(\varphi|_{\mathcal{E}})^{-1}$ at $z_{0}$ along a curve $\gamma$ in $\mathcal{D}{\setminus}\{a_{j}\}$ joining $z_{0}$ to $z$, as described previously. We set $\mathbf{h}(a_{j})=x_{j}.$ By the Monodromy Theorem, $\mathbf{h}(z)$ does not depend on the initial choice of $z_{0}$ and $\mathbf{h}(V)=\mathcal{E}.$ The set $\mathcal{C}=\mathbf{h}(\mathcal{D})$ – which is constructed using analytic continuations along curves – is naturally path connected and satisfies
@@
In order to show that $\overline{\mathcal{C}}\subset\mathbb{D}$, assume that $\mathcal{D}=g(D)$, where $g$ is a univalent map defined on a region containing the closure of a round disk $D.$ For a small $\epsilon>0$, the $\epsilon$-neighborhood of $D$ is contained in the domain of $g.$ Let $\mathcal{D}_{\epsilon}=g(D_{\epsilon}).$ Using the same process in the construction of $\mathcal{C},$ we construct another region $\mathcal{C}_{1}$ containing the closure of $\mathcal{C}$ by considering all possible analytic continuations of germs of $(\varphi|_{\mathcal{E}})^{-1}$ along curves in $\mathcal{D}_{\epsilon}$. It should be noticed, however, that now $\mathcal{D}_{\epsilon}$ might contain a finite set of ramified points, and as a consequence, some analytic continuations do not lead to a germ, but rather to a branch point of $\varphi$ which we add to $\mathcal{C}_{1}$ in the process of its construction. It follows that $\overline{\mathcal{C}}\subset\mathcal{C}_{1}\subset\mathbb{D}.$ If $K$ is a compact subset of $\mathcal{D}$, then by (6.1) the intersection $\varphi^{-1}(K)\cap\partial\mathcal{C}$ is empty. Hence
@@
is compact. Since $K$ is arbitrary, $\varphi:\mathcal{C}\to\mathcal{D}$ is proper, with the same degree $d$ of $\varphi|_{\mathcal{E}}.$ Finally, we will prove that $\mathcal{C}=\tilde{\mathcal{D}}_{j}.$ Since $\mathcal{C}$ is connected and $x_{j}\in\mathcal{C}$, it follows that $\mathcal{C}\subset\tilde{\mathcal{D}}_{j}.$ For the other inclusion, we consider a curve $\zeta$ in $\tilde{\mathcal{D}}_{j}$ joining some $w_{0}\in\mathcal{E}$ to an arbitrary $w$ in $\tilde{\mathcal{D}}_{j}.$ Without loss of generality, we may assume that $\gamma=\varphi\circ\zeta$ contains no ramified points; by Theorem 3.2, $\zeta$ is given by an analytic continuation of a germ of $(\varphi|_{\mathcal{E}})^{-1}$ along $\gamma$, thereby showing that $w$ belongs to $\mathcal{C}.$
@@
for every $G$ in $\mathcal{G}_{1}$ and every $w\in{\tilde{\mathcal{D}}}_{1}.$ In Diagram (6.2), $\varphi$ is an isometry on each disk (using the hyperbolic metric $\mu$ on the unit disk and the orbifold metric $\rho$ on $\mathbb{C}{\setminus}R$). We conclude that $\|g^{\prime}(z)\|_{\rho}<\eta_{1},$ whenever $z\in V_{1}$, for every univalent branch $g$ defined on $V_{1}.$ We may take $V_{1}$ as the open ball $B(\zeta_{0},r_{0})$ described in the statement. The Lemma follows in the case $a_{j}=a.$ The general case involves a similar reasoning (using another sequence of univalent maps just like Diagram (6.2), with the help of Lemma 6.1).
@@
Follows from Theorem B and Lemma 6.2, using a standard compactness argument, with a finite covering by relatively compact open sets on which all (finitely many) branches of $\mathbf{f}_{a}^{-1}$ contract the orbifold metric by a uniform factor. Lemma 6.2 is applied to find this uniform factor on a neighborhood of every ramified point. Using the polar representation of $\mathbf{f}_{a}^{-1}$ and the fact that $p/q>1$ it is possible to show that $g$ extends to a univalent branch on $B(\zeta_{0},r_{0})$ by computing $g$ explicitly using polar coordinates. ∎
7. Subhyperbolicity
If $a$ is a Misiurewicz point, then by definition the critical point has a unique bounded forward orbit, the preperiodic critical orbit of $\mathbf{f}_{a}$. This critical orbit is precisely the set of ramified points which are in $K_{a}$. (Recall that every ramified point is a singularity of the orbifold metric). Since both $K_{a}$ and $\mathbb{C}{\setminus}R$ are backward invariant, it follows that $K_{a}{\setminus}R$ is also backward invariant under $\mathbf{f}_{a}.$ The following result guarantees the uniform expansion of the orbifold metric within a neighborhood of $K_{a}.$
Theorem 7.1 (Subhyperbolicity).
Suppose that $a$ is a Misiurewicz point for the family (2.1). Then $\mathbf{f}_{a}$ expands the orbifold metric $\rho$ on a neighborhood of $K_{a}$ by a uniform factor. More precisely, there exist an open set $V$ containing $K_{a}$ and a constant $\eta\in(0,1)$ such that $R\cap V$ is the preperiodic critical orbit of $\mathbf{f}_{a}$, and for every univalent branch $g$ of $\mathbf{f}_{a}^{-1}$ defined on a region $W\subset V,$ we have $\|g^{\prime}(w)\|_{\rho}<\eta$, for every $w\in W{\setminus}R.$
@@ 7. Subhyperbolicity
Since the set $R$ is locally finite and $K_{a}$ is compact, there exists an $\epsilon$-neighborhood $V=(K_{a})_{\epsilon}$ such that $R\cap V=R\cap K_{a}.$
@@
For every ramified point $a_{j}$ of the canonical orbifold associated to a Misiurewicz point there exists a coordinate chart $z=\phi(\zeta)$ defined for $\zeta$ in a neighborhood of $a_{j}$, with $\phi(a_{j})=0$, such that the expression of the orbifold metric with respect to the local uniformizing parameter $z$ around zero becomes
@@
where $d=\nu(a_{j})$ and $\rho_{j}$ is a $C^{\infty}$ and strictly positive function defined on a round disk centered at 0 which satisfies $\rho_{j}(e^{2\pi i/d}z)=\rho_{j}(z),$ for any $z$ in the domain of $\rho_{j}.$
@@
The statement can be generalized for any regular branched covering $\varphi$ (not necessarily the one used in this paper). Let $x_{j}$ be a branch point. Up to compositions with translations, $\varphi$ is locally conjugate to $z^{\nu(a_{j})}$, where $a_{j}=\varphi(x_{j})$, as a consequence of Böttcher’s Theorem. The expression (7.1) is found in [12, p. 211], using the branched covering $z^{\nu(a_{j})}$ around zero, computing directly the push-forward of the metric. ∎
Recall that in the case of a Misiurewicz point, the preperiodic critical orbit of $\mathbf{f}_{a}$ contains a unique cycle $\alpha(a).$
@@
Since every cycle is a bounded orbit, it lies in the filled Julia set. The points of $\alpha(a)$ are in $R$ and none of them is the critical point. Denote the points of $\alpha(a)$ by
@@
There exist univalent branches $f_{i}$ of $\mathbf{f}_{a}$, each of which is locally defined at $\zeta_{i-1}$, satisfying $f_{i}(\zeta_{i-1})=\zeta_{i}$. Consider the composition
@@
Let $K=K_{a}.$ Recall that the points of the preperiodic orbit of zero coincide with $R\cap K,$ where $R$ is the set of ramified points. Let $R^{*}=R{\setminus}\{0\}.$ We know that every point $z$ of $K\cap R^{*}$ has only one image in $K$ (otherwise there would be another bounded forward orbit of zero), and the other q1q-1 images of $z$ are outside the compact set $K.$ By continuity, the correspondence maps a small conformal disk containing $z\in K\cap R^{*}$ to $q$ disjoint conformal disks, and only one of them intersects $K.$ This determines, at each such $z$, a special univalent branch $g_{z}$ of the correspondence that sends a conformal disk containing $z$ to the aforementioned conformal disk intersecting $K.$ The disjoint union of such small disks determines a neighborhood $V_{\alpha}$ of $K\cap R^{*}$ such that, if the correspondence sends a point $\zeta\in V_{\alpha}$ to another point $\zeta^{\prime}$ in $V_{\alpha}$, then $\zeta^{\prime}=g_{z}(\zeta)$, for some $z\in K\cap R^{*}.$ This creates on $V_{\alpha}$ a special regime of iteration; and since $\alpha(a)$ is repelling, any infinite forward orbit contained in $V_{\alpha}$ is preperiodic and eventually coincides with $\alpha(a).$
@@
For any conformal disk $U$ intersecting $K_{a}$, there exists a finite sequence of nonconstant holomorphic maps $(f_{j})_{j=0}^{n}$, defined on $U$ and starting with the identity map $f_{0}=\mathrm{Id}$, such that each $f_{j}$ is a branch of the composition $\mathbf{f}_{a}\circ f_{j-1}$, $f_{n}(U)$ contains the origin, and $f_{j}^{\prime}(z)\neq 0$ for all $z\in U$ and all $j\in\{0,\dots,n\}$. Since $U$ is arbitrary, (7.4) follows.
@@
Fix a bounded forward orbit $(z_{i})_{0}^{\infty}$ of a point in $U\cap K_{a}.$ If $0\in U$ or some point of this orbit is a critical point, then there is nothing to prove. Otherwise, by the Monodromy Theorem and its Corollary 3.1 there exists a unique branch $f_{1}$ of $\mathbf{f}_{a}$ defined on $U$ which sends $z_{0}$ to $z_{1}$. If $f_{1}(U)$ contains 0, then there is nothing to prove. Otherwise, we may consider the holomorphic multifunction $\mathbf{f}_{a}\circ f_{1}$ defined on $U$, and using analytic continuation as before, we find a unique branch $f_{2}$ of $\mathbf{f}_{a}\circ f_{1}$ defined on $U$ which sends $z_{0}$ to $z_{2}.$ Unless $f_{2}(U)$ contains the critical point, $f_{3}$ is by definition the unique branch of $\mathbf{f}_{a}\circ f_{2}$ defined on $U$ which sends $z_{0}$ to $z_{3}.$ If we are fortunate, this argument terminates at a branch $f_{n}$ of $\mathbf{f}_{a}^{n}$ defined over $U$ such that $f_{n}(U)$ contains zero, and there is nothing else to prove. Otherwise, we find a sequence of maps $f_{n}:U\to\mathbb{C}{\setminus}\{0,a\}$, which is normal, by Montel’s Theorem. However, we will see that $f_{n}$ can never be a normal family. Suppose first that the set of subsequential limits of $(z_{i})_{0}^{\infty}$ is contained in $R\cap K.$ Then all but finitely many $z_{n}$ belong the the neighborhood $V_{\alpha}$ of $K\cap R^{*}$. Hence $(z_{i})_{0}^{\infty}$ eventually coincides with the repelling cycle $\alpha,$ which implies $f_{n}^{\prime}(z_{0})\to\infty$, as a consequence of the chain rule. We conclude from the Weierstrass convergence theorem that $f_{n}$ is not normal. If the set of subsequential limits of the bounded orbit $(z_{i})_{0}^{\infty}$ is not contained in $K\cap R,$ then a subsequence $z_{n_{k}}$ converges to some $w_{0}$ in $K{\setminus}R.$ Suppose for a moment that $f_{n}$ is a normal family. Since $f_{n_{k}}(z_{0})$ converges to $w_{0},$ no subsequence of $f_{n_{k}}$ escapes to infinity. By normality, we may replace $f_{n_{k}}$ by one of its convergent subsequences, so that $f_{n_{k}}$ converges locally uniformly to some holomorphic function $g$ defined on $U$, as well as $f^{\prime}_{n_{k}}$ converges locally uniformly to $g^{\prime}.$ Since $R$ is forward invariant and $K\cap R$ is the unique bounded forward orbit of zero, it is possible to show that if one point of the sequence $(z_{i})_{0}^{\infty}$ enters $R$, then subsequent terms never leave $R$ and are eventually trapped in the repelling cycle $\alpha(a)$, a possibility that is ruled out by the fact $z_{n_{k}}\to w_{0}\notin R.$ Hence we are allowed to evaluate the norm of $f_{n_{k}}^{\prime}$ with respect to the orbifold metric $\rho$, concluding from Theorem 7.1 and the chain rule that $\|f_{n_{k}}^{\prime}\|_{\rho}$ explodes to infinity as $k\to\infty$; nevertheless, we also have that
@@
As described in (7.2) and (7.3), the composition of the univalent branches along the points $\zeta_{j}$ of the repelling cycle $\alpha$ yields a holomorphic map $f$ defined on a neighborhood of $\zeta_{0}$ with a repelling fixed point at $\zeta_{0}=f(\zeta_{0}).$ The inverse $f^{-1}$ maps a conformal disk $\Omega$ around $\zeta_{0}$ into itself, with a geometrically attracting fixed point of multiplier $\lambda_{0}$ at $\zeta_{0}.$ We may choose $\Omega$ sufficiently small so that $\zeta_{0}$ is the unique ramified point contained in $\Omega$. Let $\mathbb{D}_{r}$ denote the open disk of radius $r$ centered at zero. Using the Königs linearization theorem, we may assume that $\Omega=\varphi(\mathbb{D}_{r}),$ where $\varphi$ is a conformal conjugacy between $z\mapsto\lambda_{0}z$ and $f^{-1}$.